At T0 rod A and rod B have the same length L0. However, when they are heated to a final temperature Tf, rod A is longer than rod B by an amount d. Derive an expression for the difference in the respective coefficients of linear expansion of the two rods, αA−αB.

Express your answer in terms of the variables T0, Tf, L0 and d.
Answer #1

Coefficient of linear expansion of rod A, LA - Lo QA = LT - TO)

LA = Lo(1+QAT: – T.))

Coefficient of linear expansion of rod B, LB - Lo AB - L.T. -TO)

LB = Lo(1+ab(T- To))

Given that, p=&1-17

that is Lo(1+a[T -T%)) - Lo(1+aBT - T6)) = 0

LoQAT: – T.) - Load(T. -T.) = d

CA - B = LT - TᎧ

Difference coefficients of linear expansion of two rods is CA - B = LT - TᎧ